# A posteriori error estimation for an interior penalty virtual element method of Kirchhoff plates

### v2

@@ 3.2   The IPVEM with inhomogeneous boundary value conditions
$S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w)=h_{K}^{-2}\sum\limits_{i=1}^{n_{K}}\chi_{i}(v-\Pi_{K}^{\Delta}v)\chi_{i}(w-\Pi_{K}^{\Delta}w),$
where $\{\chi_{i}\}$ are the local DoFs on $K$ with $n_{K}$ being the number of the DoFs. Let
$J_{1}(v,w)=\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[v][w]\mathrm{d}s+\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$ (3.9)
with $\lambda_{e}\geq 1$ being some edge-dependent parameter, and the additional two terms are given by
$J_{2}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}w}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
Here and bellows, we define the piecewise $H^{2}$ projector $\Pi_{h}^{\Delta}$ by setting $\Pi_{h}^{\Delta}|_{K}=\Pi_{K}^{\Delta}$ for all $K\in\mathcal{T}_{h}$.
The IPVEM for (2.1) with inhomogeneous boundary value conditioins can be described as follows: Find $u_{h}\in V_{h}^{g}=\{v\in V_{h}:v=I_{h}g_{D}~{}{\rm on}~{}\partial\Omega\}$ such that
@@
where $g_{D}$ has been extended to the whole domain. The bilinear form is
where $a_{h}(v,w)=\sum_{K\in\mathcal{T}_{h}}a_{h}^{K}(v,w)$. The right-hand side is
$F_{h}(v_{h})=\langle f_{h},v_{h}\rangle+\sum_{e\in\mathcal{E}_{h}^{\partial}}\int_{e}\Big{(}-\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}+\frac{\lambda_{e}^{3}}{|e|^{3}}[u][v_{h}]+\frac{\lambda_{e}}{|e|}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{)}\mathrm{d}s,$
@@
with $\Pi_{0,K}^{k}$ being the $L^{2}$ projector onto $\mathbb{P}_{k}(K)$. Then for $f\in H^{k-2}(\Omega)$ with $k\geq 2$, we have
@@
With the mesh-dependent parameter $\lambda_{e}$ chosen as in [18], we can deduce the stability result by using the same arguments. We omit the details with the results described as follows.
@@
For later use, we provide the following norm equivalence with the proof given in [29].
@@
For $m=0,1,2$, there holds
$|v-\Pi_{K}^{\Delta}v|_{m,K}\eqsim h_{K}^{1-m}\|\bm{\chi}(v-\Pi_{K}^{\Delta}v)\|_{\ell^{2}},\qquad v\in V_{k}(K).$ (3.17)
@@ 4 Enriching operator
This section focuses on proposing and analyzing an enriching operator $E_{h}$ from $V_{h}$ to $H^{2}$-conforming virtual element spaces. The enriching operator is a special kind of quasi-interpolation operator, which connects a finite dimensional space to another one with higher order regularity. We refer to [5, 9, 19, 22, 20, 8] and the references therein for the construction and application of such operators in finite element methods. Such an operator in the virtual element context was first proposed in [21] and applied in the medius error analysis of some nonconforming VEMs. This idea was also adopted in [16] for the a posteriori error estimation for a $C^{1}$ virtual element method of Kirchhoff plates.
@@
Note that the above moments vanish in the lowest order case $k=2$.
@@
Let ${\mathcal{E}_{a}^{0}}$ and $\mathcal{E}_{e}^{0}$ denote the interior vertices and edges of $\mathcal{T}_{h}$, respectively. The global dual space is defined as
@@
    $\mathcal{N}_{a}^{z}(v)=v(z),\quad z\in\mathcal{E}_{a}^{0}.$
@@
    $\mathcal{N}_{\nabla}^{z}(v)=h_{z}\nabla v(z),\quad z\in\mathcal{E}_{a}^{0}.$
@@
    $\mathcal{N}_{e}(v)=|e|^{-1}(m_{e},v)_{e},\quad m_{e}\in\mathbb{M}_{k-4}(e),\quad e\in\mathcal{E}_{e}^{0}.$
@@
    $\mathcal{N}_{n_{e}}(v)=(m_{e},\partial_{\bm{n}}v)_{e},\quad m_{e}\in\mathbb{M}_{k-3}(e),\quad e\in\mathcal{E}_{e}^{0}.$
@@
    $\mathcal{N}_{a}^{z}(E_{h}\phi)=(E_{h}\phi)(z):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}\Pi_{h}^{\Delta}\phi|_{K^{\prime}}(z).$
@@
    $\mathcal{N}_{\nabla}^{z}(E_{h}\phi):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}h_{z}\nabla\Pi_{h}^{\Delta}\phi|_{K^{\prime}}(z).$
@@
    $\mathcal{N}_{e}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}|e|^{-1}\int_{e}m_{e}\Pi_{h}^{\Delta}\phi|_{K^{\prime}}\mathrm{d}s,\quad m_{e}\in\mathbb{M}_{k-4}(e).$
@@
    $\mathcal{N}_{n_{e}}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}\int_{e}m_{e}\partial_{\bm{n}}\Pi_{h}^{\Delta}\phi|_{K^{\prime}}{\text{d}}s,\quad m_{e}\in\mathbb{M}_{k-3}(e).$
@@
    $\mathcal{N}_{K}(E_{h}\phi):=|K|^{-1}\int_{K}m_{K}\Pi_{h}^{\Delta}\phi|_{K}{\mathrm{d}}x,\quad m_{K}\in\mathbb{M}_{k-4}(K).$
@@
$|v-E_{h}v|_{2,h}^{2}\lesssim\sum_{K\in\mathcal{T}_{h}}S^{K}(v-\Pi_{h}^{\Delta}v,v-\Pi_{h}^{\Delta}v)+\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)},$ (4.2)
@@
According to the triangle inequality
$|v-E_{h}v|_{2,h}\leq|v-\Pi_{h}^{\Delta}v|_{2,h}+|\Pi_{h}^{\Delta}v-E_{h}v|_{2,h}$
and the norm equivalence (3.17), it reduces to bound the second term on the right-hand side. Since $\Pi_{h}^{\Delta}v-E_{h}v\in V_{k}^{2,c}(K)$, by the inverse inequality and norm equivalence in [21], one gets
$|\Pi_{h}^{\Delta}v-E_{h}v|_{2,K}\lesssim h_{K}^{-2}\|\Pi_{h}^{\Delta}v-E_{h}v\|_{0,K}\lesssim h_{K}^{-1}\|\bm{\chi}^{c}(\Pi_{h}^{\Delta}v-E_{h}v)\|_{l^{2}},\qquad K\in\mathcal{T}_{h}.$
@@
(1) Let $z\in\mathcal{V}_{h}^{0}$ be an interior vertex. Assume that $K_{1}=K,K_{2},\cdots,K_{L}$ share the node $z$, and denote $(\Pi_{h}^{\Delta}\phi)_{i}=\Pi_{h}^{\Delta}\phi|_{K_{i}}$, where $K_{i}$ and $K_{i+1}$ are two neighboring elements. We have
$(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)$
$=(\Pi_{h}^{\Delta}v)_{1}(z)-\frac{1}{L}((\Pi_{h}^{\Delta}v)_{1}+\cdots+(\Pi_{h}^{\Delta}v)_{L})(z)$
$=\frac{1}{L}(((\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2})+((\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{3})+\cdots+((\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{L}))(z).$
@@
$(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{j}(z)=\sum\limits_{i=1}^{j-1}((\Pi_{h}^{\Delta}v)_{i}-(\Pi_{h}^{\Delta}v)_{i+1})(z),$
it suffices to consider the term $(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)$. Apply the inverse inequality of polynomials to get
$|(\Pi_{h}^{\Delta}v)_{1}(z)-(\Pi_{h}^{\Delta}v)_{2}(z)|\leq\|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}\|_{\infty,e}\lesssim h_{e}^{-1/2}\|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}.$ (4.3)
@@
$\|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}$ $=\|(\Pi_{h}^{\Delta}v)_{1}-v_{1}+v_{1}-v_{2}+v_{2}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}$
$\leq\|(\Pi_{h}^{\Delta}v)_{1}-v_{1}\|_{0,e}+\|v_{2}-(\Pi_{h}^{\Delta}v)_{2}\|_{0,e}+\|[v]\|_{0,e}.$
@@
$\|v-\Pi_{h}^{\Delta}v\|_{0,e}\lesssim h_{K}^{1/2}|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1/2}\|v-\Pi_{h}^{\Delta}v\|_{0,K}.$
@@
$|(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)|=|\Pi_{h}^{\Delta}v(z)-v(z)|\leq\|\Pi_{h}^{\Delta}v-v\|_{\infty,K}.$
By the Sobolev inequality (2.6),
$\|v-\Pi_{h}^{\Delta}v\|_{\infty,K}\leq h_{K}|v-\Pi_{h}^{\Delta}v|_{2,K}+|v-\Pi_{h}^{\Delta}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\Delta}v\|_{0,K}.$
Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using the norm equivalence (3.17), we obtain
$\Big{(}h_{K}^{-1}(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim S^{K}(v-\Pi_{h}^{\Delta}v,v-\Pi_{h}^{\Delta}v)+\sum_{e_{i}}\frac{1}{|e_{i}|^{3}}\|[v]\|_{0,e_{i}}^{2},$
@@
(2) For the gradient value at an interior vertex $z\in\mathcal{V}^{0}$, one has
$h_{z}\nabla(\Pi_{h}^{\Delta}v-E_{h}v)|_{K}(z)$
$=$ $h_{z}(\nabla\Pi_{h}^{\Delta}v)_{1}(z)-\frac{h_{z}}{L}((\nabla\Pi_{h}^{\Delta}v)_{1}+\cdots+(\nabla\Pi_{h}^{\Delta}v)_{L})(z)$
$=$ $\frac{h_{z}}{L}(((\nabla\Pi_{h}^{\Delta}v)_{1}-(\nabla\Pi_{h}^{\Delta}v)_{2})+((\nabla\Pi_{h}^{\Delta}v)_{1}-(\nabla\Pi_{h}^{\Delta}v)_{3})+\cdots+((\nabla\Pi_{h}^{\Delta}v)_{1}-(\nabla\Pi_{h}^{\Delta}v)_{L}))(z)$
and we only need to consider the term $h_{z}((\nabla\Pi_{h}^{\Delta}v)_{1}(z)-(\nabla\Pi_{h}^{\Delta}v)_{2}(z))$. By the inverse inequality of polynomials,
$h_{z}((\nabla\Pi_{h}^{\Delta}v)_{1}(z)-(\nabla\Pi_{h}^{\Delta}v)_{2}(z))$ $\lesssim h_{z}h_{e}^{-1/2}|(\Pi_{h}^{\Delta}v)_{1}-(\Pi_{h}^{\Delta}v)_{2}|_{1,e}$
$\lesssim h_{e}^{1/2}\Big{(}|(\Pi_{h}^{\Delta}v)_{1}-v_{1}|_{1,e}+|v_{2}-(\Pi_{h}^{\Delta}v)_{2}|_{1,e}+|[v]|_{1,e}\Big{)}.$
Using the trace inequality, we get
$|v-\Pi_{h}^{\Delta}v|_{1,e}\lesssim h_{K}^{1/2}|v-\Pi_{h}^{\Delta}v|_{2,K}+h_{K}^{-1/2}|v-\Pi_{h}^{\Delta}v|_{1,K}.$

### v4

@@ 3.2   The discrete IP bilinear form
$S^{K}(v,w)=\sum\limits_{i=1}^{n_{K}}\chi_{i}(v)\chi_{i}(w),$
where $\{\chi_{i}\}$ are the local DoFs on $K$ with $n_{K}$ being the number of the DoFs. As was done in [18], we define
$J_{1}(v,w)=\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$ (3.10)
$J_{2}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
Here and below, we define the piecewise $H^{1}$ and $H^{2}$ projectors $\Pi_{h}^{\nabla}$ and $\Pi_{h}^{\Delta}$ by setting $\Pi_{h}^{\nabla}|_{K}=\Pi_{K}^{\nabla}$ and $\Pi_{h}^{\Delta}|_{K}=\Pi_{K}^{\Delta}$ for all $K\in\mathcal{T}_{h}$.
The bilinear form is
where $a_{h}(v,w)=\sum_{K\in\mathcal{T}_{h}}a_{h}^{K}(v,w)$. The discrete scheme of IPVEM solving problem (2.1) is to find $u_{h}\in V_{h}$ such that
@@
The right-hand side is
$F_{h}(v_{h}):=\langle f_{h},v_{h}\rangle,$
@@
with $\Pi_{0,K}^{k}$ being the $L^{2}$ projector onto $\mathbb{P}_{k}(K)$. For $f\in H^{k-2}(\Omega)$ with $k\geq 2$, we have
@@
Remark 3.1.
As observed in [18], despite the discontinuity, the interior penalty virtual elements display behaviors similar to $C^{0}$-continuous elements in the a priori estimate. We will demonstrate that this observation also holds for the a posteriori estimate. Therefore, we do not include the penalty term
$\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[v][w]\mathrm{d}s$
in the discrete scheme and introduce the error estimator
$\eta_{0,K}=\Big{(}\sum\limits_{e\in\mathcal{E}(K)\cap\mathcal{E}_{h}}\eta_{0,e}^{2}\Big{)}^{1/2},\qquad\eta_{0,e}=\frac{1}{|e|^{3/2}}\|[u_{h}]\|_{0,e}$ (3.15)
in the a posteriori estimate, as will be discussed later.
@@
Referring to [18], the mesh-dependent parameter $\lambda_{e}\eqsim 1$. Still, we can deduce the stability result by using the same arguments. We omit the details with the results described as follows.
@@
For later use, we provide the following norm equivalence.
@@
For every $v\in V_{k}(K)$, there hold
$|v-\Pi_{K}^{\Delta}v|_{2,K}\eqsim h_{K}^{-1}\|\bm{\chi}(v-\Pi_{K}^{\Delta}v)\|_{\ell^{2}},$ (3.19)
$|v-\Pi_{K}^{\nabla}v|_{1,K}\eqsim\|\bm{\chi}(v-\Pi_{K}^{\nabla}v)\|_{\ell^{2}},$ (3.20)
$h_{K}^{-1}\|v-\Pi_{K}^{\nabla}v\|_{0,K}\eqsim\|\bm{\chi}(v-\Pi_{K}^{\nabla}v)\|_{\ell^{2}}.$ (3.21)
Proof.
The first and second equations are presented in [27, Lemma 4.1]. Following the approach in [21], it is straightforward to derive the lower bound estimate $\|\bm{\chi}(v-\Pi_{K}^{\nabla}v)\|_{\ell^{2}}\lesssim h_{K}^{-1}\|v-\Pi_{K}^{\nabla}v\|_{0,K}$ through direct calculations.
For the upper bound estimate, according to the boundedness of $\Pi_{h}^{\nabla}$ and $\Pi_{h}^{\Delta}$ in Lemma 3.3 and the Poincaré-Friedrichs inequality (2.7), one has
$\|v-\Pi_{h}^{\nabla}v\|_{0,K}$ $\leq\|v-\Pi_{h}^{\Delta}v\|_{0,K}+\|\Pi_{h}^{\Delta}v-\Pi_{h}^{\nabla}v\|_{0,K}$
$=\|v-\Pi_{h}^{\Delta}v\|_{0,K}+\|\Pi_{h}^{\nabla}(\Pi_{h}^{\Delta}v-v)\|_{0,K}\lesssim\|v-\Pi_{h}^{\Delta}v\|_{0,K}$
$\lesssim h_{K}|v-\Pi_{h}^{\Delta}v|_{1,K}\leq h_{K}|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}|\Pi_{h}^{\Delta}(\Pi_{h}^{\nabla}v-v)|_{1,K}$
$\lesssim h_{K}|v-\Pi_{h}^{\nabla}v|_{1,K}.$ (3.22)
This along with the second formula yields the desired result. ∎
@@ 4 Enriching operator
This section focuses on proposing and analyzing an enriching operator $E_{h}$ from $V_{h}$ to $H^{2}$-conforming virtual element spaces. The enriching operator is a special kind of quasi-interpolation operator, which connects a finite-dimensional space to another one with higher-order regularity. We refer to [5, 9, 19, 22, 20, 8] and the references therein for the construction and application of such operators in finite element methods. Such an operator in the virtual element context was first proposed in [21] and applied in the medius error analysis of some nonconforming VEMs. This idea was also adopted in [16] for the a posteriori error estimation for a $C^{1}$ virtual element method of Kirchhoff plates.
@@
Note that the above moments vanish in the lowest-order case $k=2$.
@@
The global dual space is defined as
@@
    $\mathcal{N}_{a}^{z}(v)=v(z),\quad z\in\mathcal{V}_{h}^{0}.$
@@
    $\mathcal{N}_{\nabla}^{z}(v)=h_{z}\nabla v(z),\quad z\in\mathcal{V}_{h}^{0}.$
@@
    $\mathcal{N}_{e}(v)=|e|^{-1}(m_{e},v)_{e},\quad m_{e}\in\mathbb{M}_{k-4}(e),\quad e\in\mathcal{E}_{h}^{0}.$
@@
    $\mathcal{N}_{n_{e}}(v)=(m_{e},\partial_{\bm{n}}v)_{e},\quad m_{e}\in\mathbb{M}_{k-3}(e),\quad e\in\mathcal{E}_{h}^{0}.$
@@
    $\mathcal{N}_{a}^{z}(E_{h}\phi)=(E_{h}\phi)(z):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}\Pi_{h}^{\nabla}\phi|_{K^{\prime}}(z)\quad z\in\mathcal{V}_{h}^{0}.$
@@
    $\mathcal{N}_{\nabla}^{z}(E_{h}\phi):=\frac{1}{N(z)}\sum\limits_{K^{\prime}\in\omega(z)}h_{z}\nabla\Pi_{h}^{\nabla}\phi|_{K^{\prime}}(z)\quad z\in\mathcal{V}_{h}^{0}.$
@@
    $\mathcal{N}_{e}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}|e|^{-1}\int_{e}m_{e}\Pi_{h}^{\nabla}\phi|_{K^{\prime}}\mathrm{d}s,\quad m_{e}\in\mathbb{M}_{k-4}(e).$
@@
    $\mathcal{N}_{n_{e}}(E_{h}\phi):=\frac{1}{N(e)}\sum\limits_{K^{\prime}\in\omega(e)}\int_{e}m_{e}\partial_{\bm{n}}\Pi_{h}^{\nabla}\phi|_{K^{\prime}}{\text{d}}s,\quad m_{e}\in\mathbb{M}_{k-3}(e).$
@@
    $\mathcal{N}_{K}(E_{h}\phi):=|K|^{-1}\int_{K}m_{K}\Pi_{h}^{\nabla}\phi|_{K}{\mathrm{d}}x,\quad m_{K}\in\mathbb{M}_{k-4}(K).$
@@
$|v-E_{h}v|_{2,h}^{2}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v)+\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\Big{)}.$ (4.2)
In addition,
$\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|^{3}}\|[v]\|_{0,e}^{2}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v),$ (4.3)
$\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v)+\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}.$ (4.4)
@@
Step 1: We first prove (4.2). Consider the triangle inequality
$|v-E_{h}v|_{2,h}\leq|v-\Pi_{h}^{\nabla}v|_{2,h}+|\Pi_{h}^{\nabla}v-E_{h}v|_{2,h}.$ (4.5)
Using the inverse inequality (3.6) and the norm equivalence (3.20), we obtain
$|v-\Pi_{h}^{\nabla}v|_{2,h}$ $=\sum_{K\in\mathcal{T}_{h}}|v-\Pi_{h}^{\nabla}v|_{2,K}\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-1}|v-\Pi_{h}^{\nabla}v|_{1,K}$
$\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{-1}(S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v))^{1/2}.$
Thus, it reduces to bound the second term on the right-hand side of (4.5). Since $\Pi_{h}^{\nabla}v-E_{h}v\in V_{k}^{2,c}(K)$, by the inverse inequality and the norm equivalence in [21], one gets
$|\Pi_{h}^{\nabla}v-E_{h}v|_{2,K}\lesssim h_{K}^{-2}\|\Pi_{h}^{\nabla}v-E_{h}v\|_{0,K}\lesssim h_{K}^{-1}\|\bm{\chi}^{c}(\Pi_{h}^{\nabla}v-E_{h}v)\|_{l^{2}},\quad K\in\mathcal{T}_{h}.$
@@
Let $z\in\mathcal{V}_{h}^{0}$ be an interior vertex. Assume that $K_{1}=K,K_{2},\cdots,K_{L}$ share the node $z$, and denote $(\Pi_{h}^{\nabla}\phi)_{i}=\Pi_{h}^{\nabla}\phi|_{K_{i}}$, where $K_{i}$ and $K_{i+1}$ are two neighboring elements. We have
$(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)$
$=(\Pi_{h}^{\nabla}v)_{1}(z)-\frac{1}{L}((\Pi_{h}^{\nabla}v)_{1}+\cdots+(\Pi_{h}^{\nabla}v)_{L})(z)$
$=\frac{1}{L}(((\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2})+((\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{3})+\cdots+((\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{L}))(z).$
@@
$(\Pi_{h}^{\nabla}v)_{1}(z)-(\Pi_{h}^{\nabla}v)_{j}(z)=\sum\limits_{i=1}^{j-1}((\Pi_{h}^{\nabla}v)_{i}-(\Pi_{h}^{\nabla}v)_{i+1})(z),$
it suffices to consider the term $(\Pi_{h}^{\nabla}v)_{1}(z)-(\Pi_{h}^{\nabla}v)_{2}(z)$. Apply the inverse inequality of polynomials to get
$|(\Pi_{h}^{\nabla}v)_{1}(z)-(\Pi_{h}^{\nabla}v)_{2}(z)|\leq\|(\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2}\|_{\infty,e}\lesssim h_{e}^{-1/2}\|(\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2}\|_{0,e}.$ (4.6)
@@
$\|(\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2}\|_{0,e}$ $=\|(\Pi_{h}^{\nabla}v)_{1}-v_{1}+v_{1}-v_{2}+v_{2}-(\Pi_{h}^{\nabla}v)_{2}\|_{0,e}$
$\leq\|(\Pi_{h}^{\nabla}v)_{1}-v_{1}\|_{0,e}+\|v_{2}-(\Pi_{h}^{\nabla}v)_{2}\|_{0,e}+\|[v]\|_{0,e}.$
@@
$\|v-\Pi_{h}^{\nabla}v\|_{0,e}\lesssim h_{K}^{1/2}|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}^{-1/2}\|v-\Pi_{h}^{\nabla}v\|_{0,K}.$
@@
$|(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)|=|\Pi_{h}^{\nabla}v(z)-v(z)|\leq\|\Pi_{h}^{\nabla}v-v\|_{\infty,K}.$
Using the Sobolev inequality (2.6) and the inverse inequality (3.6), we get
$\|v-\Pi_{h}^{\nabla}v\|_{\infty,K}$ $\leq h_{K}|v-\Pi_{h}^{\nabla}v|_{2,K}+|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\nabla}v\|_{0,K}$
$\lesssim|v-\Pi_{h}^{\nabla}v|_{1,K}+h_{K}^{-1}\|v-\Pi_{h}^{\nabla}v\|_{0,K}.$
Therefore, for any vertex $z$ of $\mathcal{T}_{h}$, by using norm equivalence (3.20) and (3.21), we obtain
$\Big{(}h_{K}^{-1}(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)\Big{)}^{2}\lesssim h_{K}^{-2}S^{K}(v-\Pi_{h}^{\nabla}v,v-\Pi_{h}^{\nabla}v)+\sum_{e_{i}}\frac{1}{|e_{i}|^{3}}\|[v]\|_{0,e_{i}}^{2},$
@@
For the gradient value at an interior vertex $z\in\mathcal{V}^{0}$, one has
$h_{z}\nabla(\Pi_{h}^{\nabla}v-E_{h}v)|_{K}(z)$
$=$ $h_{z}(\nabla\Pi_{h}^{\nabla}v)_{1}(z)-\frac{h_{z}}{L}((\nabla\Pi_{h}^{\nabla}v)_{1}+\cdots+(\nabla\Pi_{h}^{\nabla}v)_{L})(z)$
$=$ $\frac{h_{z}}{L}(((\nabla\Pi_{h}^{\nabla}v)_{1}-(\nabla\Pi_{h}^{\nabla}v)_{2})+\cdots+((\nabla\Pi_{h}^{\nabla}v)_{1}-(\nabla\Pi_{h}^{\nabla}v)_{L}))(z)$
and we only need to consider the term $h_{z}((\nabla\Pi_{h}^{\nabla}v)_{1}(z)-(\nabla\Pi_{h}^{\nabla}v)_{2}(z))$. By the inverse inequality of polynomials,
$h_{z}((\nabla\Pi_{h}^{\nabla}v)_{1}(z)-(\nabla\Pi_{h}^{\nabla}v)_{2}(z))\lesssim$ $h_{z}h_{e}^{-1/2}|(\Pi_{h}^{\nabla}v)_{1}-(\Pi_{h}^{\nabla}v)_{2}|_{1,e}$
$\lesssim$ $h_{e}^{1/2}(|(\Pi_{h}^{\nabla}v)_{1}-v_{1}|_{1,e}+|v_{2}-(\Pi_{h}^{\nabla}v)_{2}|_{1,e}+|[v]|_{1,e}).$
Using the trace inequality (2.5) and inverse inequality (3.6), we get
$|v-\Pi_{h}^{\nabla}v|_{1,e}\lesssim h_{K}^{-1/2}|v-\Pi_{h}^{\nabla}v|_{1,K}.$